section 8

Kosovski Company is considering Projects S and L, whose cash flows are shown below. These projects are
mutually exclusive, equally risky, and are not repeatable. If the decision is made by choosing the project with
the higher IRR, how much value will be forgone? Note that under some conditions choosing projects on the
basis of the IRR will cause $0.00 value to be lost.
WACC: 7.75%
Year 0 1 2 3 4
CFS ?$1,050 $675 $650
CFL ?$1,050 $360 $360 $360 $360
____
Answer #1
Calculating the Net present value of each project:
From the above table, column-4 reprsents the WACC at 7.75% calculated using he below formula.
Factor = 1 / (1+r)^n
where r is the discount rate = 7.75%
           n is the number of that particular period.
For year-1: 1 / (1+0.0775)^1 = 0.9281
For year-2: 1 / (1+0.0775)^2 = 0.8613
Similarly we have to calculate for the remaining years.
The present value of cash flows is obtained by multiplying the cash flows of each project with the discount factor of that particular period. This process is done till year-4. After calculating the present value of cash flows, we must add those from year-0 to year-4 to get the NPV of each project.
Hence the NPV of the Project-S is $136.31 and the NPV of project-L is $149.03. The project with the highest NPV is choosen as the best project. Since the projects are mutually exclusive project-L is choosen as the best one based on the NPV criterion.
If the decision criterion is made based on the IRR , then the values comes to
The IRR of each project is calculated using the excel sheet:
Step1: Enter the cash flows from year-0 to year-4 in a column.
Step2: Select any cell beside the column and enter the formula as "=IRR(select cash flows from year-0 to year-4)"
Step3: Enter the formula as above and click "OK" to get the desired value.
Step4: Adjust the value upto "2" decimal points.
Therefore, the IRR for project-S is 17.13% and for project-L it is 13.95%
Since the projects are mutually exclusive project-S is choosen as the best one based on the IRR criterion.